What is the length and width of 48sq ft ?

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

there are many possible answers
including
48 by 1
24 by 2
12 by 4
6 by 8
3 by 16

Answer by Alan3354(69443)

What is the length and width of 48sq ft ?
--------------
L*W = 48
L = 48/W
Pick any value for W > 0 and find L
----
W = 0.01
L = 4800
-----------
W = pi
L = 48/pi
etc.